Copper-production increased at a rate of about 4.9% per year between 1988 and 1993. In 1993, copper-production was approximately 1.801 billion kilograms.
If this trend continued, what equation will best model the copper-production (P), in billions of kilograms, since 1993.
(Let t = 0 for 1993.)

Answer by jim_thompson5910(35256)


Answer: P = 1.801*(1.049)^t

================================================
Explanation:

The general format is
P = a*(1+r)^t
where
P = final amount after t years
a = starting amount
r = growth rate in decimal form
t = number of years that have elapsed

In this case,
P = unknown
a = 1.801 (value in billions)
r = 0.049 (representing 4.9% since 4.9% = 4.9/100 = 0.049)
t = also unknown

So we go from
P = a*(1+r)^t
to
P = 1.801*(1+0.049)^t
which simplifies to
P = 1.801*(1.049)^t

This equation only works if we assume the growth rate of 4.9% holds the same for years after 1993.